Gina is going to a party and was asked to bring 3lbs of mixed nuts that includes cashews and peanuts. If the cashews cost $6.00 per pound and the peanuts cost $1.50 per pound, how many pounds of each did she buy if her total cost was $9.00?
I know the answer from doing it in my head, I just need help setting up the equation properly on paper. Thanks for your input / time!

Answer by ptaylor(2198)


Let x=number of lbs cashews
Then 3-x=number of lbs of peanuts
We know that $6.00x=cost of the cashews
We also know that $1.50(3-x)=cost of peanuts
Now we are told that her total cost was $9.00, so our equation to solve is:
$6.00x+$1.50(3-x)=$9.00 get rid of parens
$6.00x+$4.50-$1.50x=$9.00 subtract $4.50 from both sides
$6.00x+$4.40-$4.50-$1.50x=$9.00-$4.50 collect like terms
$4.50x=$4.50 divide both sides by $4.50
x=1 lb---------------------------------number of lbs of cashews
3-x=3-1=2 lbs-----------------------number of lbs of peanuts
CK
$6.00(1)+$1.50(2)=$9.00
$6.00+$3.00=$9.00
$9.00=$9.00

Hope this helps---ptaylor